# Controllable Learning in Information Retrieval: A Survey

### v1

@@ 4.1.4 Other Methods
There are also some control methods in machine learning, which is not strictly belong to the above three classifications. Multi-VAE [63] provide the ‘knobs’ in a Disentanglement (Disent.) way for users, with each knob corresponding to an item aspect. The latent space where generative factors (here, a preference towards an item category like genre) are captured independently in their respective dimensions, thereby enabling predictable manipulations. CGIR [64] propose a weakly-supervised method that can learn a disentangled item representation from user-item interaction data and ground the semantic meaning of attributes to dimensions of the item representation. During inference, CGIR start from the reference item and “walk” along the direction of the modification in the item representation space to retrieve a sequence of items in a gradient manner. Test-Time Adaption (TTA) means the algorithms do not change the model through retraining manner. especially in the scenario of online recommendation, where time consuming is quite strict, some works focus on how to change the model to adapt the dynamic environment directly in test stage. Reinforcement Learning (RL) is a kind of classic method to achieve controllability. Since these algorithms learn from the interaction with the environment (actually the reward function of the environment), it is possible for us to control the algorithms through dedicatedly designing the reward function. Inspired by the exceptional general intelligence of Large Language Models (LLMs) [48], researchers have begun to explore their application in recommender systems and other fields. [65] construct Supervised Fine-Tuning (SFT) tasks, augmented with labels derived from a conventional recommender model, explicitly improving LLMs’ proficiency in adhering to recommendation specific instruction which could be regarded as task description $s_{desc}$ as defined in Sec.2.
@@
Table I: Summary of representative controllable learning (CL) methods for information retrieval. “Who” denotes the entity responsible for control (see Section 3.1), including User-Centric Control (User) and Platform-Mediated Control (Platform). “What” refers to what is task target (see Section 3.2) including, Multi-Objective Control (MOC), Historical Behavior Control (HBC), Controllable Environmental Adaptation (CEA), Multi-Scenario Control (MSC), and Multi-Domain Control (MDC). “CL Tech.” describes the techniques for implementing control (see Section 4.1) including Rule-Based Techniques (RBT), Pareto Optimization (Pareto), Hypernetwork (HN), Disentanglement Latent Space (Disent.), Test- Time Adaption (TTA), Reinforcement Learning (RL), Supervised Fine- Tuning (SFT). “Where” indicates the position where control is applied in the testing stage (see Section 4.2) including Pre-Processing Methods (Pre.), In-Processing Methods (In.), and Post-Processing (Post.).
@@
Method Year Who What CL Tech. Where
ComiRec [5] 2020 Platform MOC RBT Post.
Mult-VAE [63] 2021 User MOC Disent. In.
CGIR [64] 2021 User HBC Disent. Pre.
UCRS [1] 2022 User MOC RBT Pre.&Post.
LACE [3] 2023 User HBC Disent. Pre.
CMR [6] 2023 Platform MOC HN In.
SAMD [57] 2023 Platform MSC HN In.
DTRN [58] 2023 Platform MSC HN In.
BFB [8] 2023 Platform MOC RL In.
HyperBandit [9] 2023 Platform CEA HN&TDA In.
HAMUR [24] 2023 Platform MDC HN In.
PHN-HVI [16] 2023 Platform MOC HN&Pareto In.
CCDF [7] 2024 Platform MOC HN In.
IFRQE [2] 2024 User HBC TDA Pre.
RecLM-gen [65] 2024 User HBC SFT&RL In.
@@ 4.2 Where to Control in Information Retrieval Models
In this section, we categorize control learning methods into three distinct categories based on the inference process during the testing phase. Specifically, we classify them according to the time when the control function defined in Definition 1 affects the process: methods that adjust model inputs before inference are termed pre-processing methods, those that operate on model parameters during inference are termed in-processing methods, and those that act on model outputs after inference are termed post-processing methods.
@@ 4.2.1 Pre-Processing Methods
Pre-processing methods, also referred to as embedding methods [27], achieve the task target solely by transforming the model inputs without adjusting the model itself, given a task description. Transformation of model inputs includes directly concatenating the task description onto the original feature vector. These methods are akin to in-context learning methods in large language models, where the task description serves as the prompt.
@@ 4.2.2 In-Processing Methods
In-processing methods adaptively adjust the parameters or hyperparameters of IR models upon receiving the task description and context to achieve the task target. The CMR framework [6] harnesses the input preference vector to guide the hypernetwork, which in turn generates network parameters tailored to the desired balance of objectives, thus achieving the desired control during the testing stage. HyperBandit [9] utilizes the periodic time information to inject to a hypernetwork, modeling the relationship between user preference with corresponding time block, achieving efficient user preference adaptation during testing stage. Similarly, the CCDF [7] employs a hyperparameter, denoted as ‘$k$’, to directly manipulate the number of categories presented in the top-$k$ recommendations, allowing for precise control over the diversity of the recommended content. These methods demonstrate the utility of in-processing strategies in enhancing the adaptability and performance of recommender systems.
@@ 4.2.3 Post-Processing Methods
Post-processing methods, including reranking and result diversification, are crucial for refining the output of recommender systems. ComiRec [5] leverages a final aggregation module that balances dual objectives—accuracy and diversity—through a weighted summation approach to determine the top-k recommendations. This method ensures that the final selection of recommendations is not only precise but also diverse, catering to a broader range of user interests. Similarly, the UCRS [1] employs a weighted summation technique that balances accuracy and coverage to mitigate filter bubbles, ensuring that users are exposed to a wider array of content beyond their established preferences. These strategies highlight the importance of post-processing in delivering a nuanced and balanced set of recommendations that align with both the system’s performance goals and the user’s discovery of diverse items. Some works achieve controllability by re-ranking. MMR [66] promote diversity of the recommendation lists generated by some models via re-ranking.
@@ 5.1.1 Single-Objective Metrics
NDCG [29] evaluates the effectiveness of information retrieval systems by accounting for both the relevance and rank of retrieved documents. It applies a logarithmic discount to lower-ranked items and normalizes against an ideal ranking for consistency. This makes NDCG essential for comparing search engine and recommender system performance across different queries.
@@
Precision is a metric used to evaluate the accuracy of information retrieval systems by measuring the proportion of retrieved documents that are relevant. It is calculated as the number of relevant documents retrieved divided by the total number of documents retrieved. High precision indicates that the system retrieves mostly relevant items, minimizing the presence of irrelevant information. This metric is particularly important in scenarios where presenting irrelevant items can negatively impact user experience, such as in search engines, recommender systems, and spam filtering.
@@
Recall measures the effectiveness of information retrieval systems by determining the proportion of relevant documents that have been successfully retrieved from the total relevant documents available. It is calculated as the number of relevant documents retrieved divided by the total number of relevant documents. High recall indicates that the system retrieves most of the relevant items, which is essential in tasks where missing relevant information can have significant consequences, such as medical research or legal discovery.
@@
Hit Rate evaluates the performance of recommender systems by assessing the presence of relevant items within the top-N recommendations provided to users. It is calculated as the number of users for whom at least one relevant item is included in the top-N recommendations, divided by the total number of users. Hit Rate is particularly useful for understanding the effectiveness of recommendation algorithms in scenarios where presenting at least one relevant option can significantly impact user satisfaction, such as in e-commerce or content streaming platforms.
@@
$\alpha$-NDCG [30] extends the NDCG metric to evaluate the diversity in information retrieval systems. By incorporating a parameter $\alpha$, it penalizes redundancy and rewards the retrieval of diverse, relevant documents. This makes $\alpha$-NDCG particularly useful for tasks like web search and recommender systems, where presenting varied content is crucial.
@@
ERR-IA [31] evaluates the effectiveness of information retrieval systems by considering user intent in the evaluation process. ERR-IA extends the Expected Reciprocal Rank (ERR) metric by incorporating a probabilistic approach to user intents, ensuring that the evaluation reflects the variety of user needs. It models user satisfaction as a function of the relevance of retrieved documents and their alignment with multiple user intents, making ERR-IA particularly useful for web search and other scenarios where understanding and catering to diverse user intents is crucial.
@@ 5.1.2 Multi-Objective Optimization Metrics
According to [32], performance indicators in multiobjective optimization can be mainly divided into four groups: Cardinality indicators, Convergence indicators, Distribution and spread indicators, and Convergence and distribution indicators. However, in practical implementations, the Pareto front approximation is not obtained through evolutionary algorithms that yield a finite (discrete) set of points. For example, [5] achieves diversity control by adjusting the continuous value $\lambda$ in the aggregation module, and similarly, [6] controls diversity and accuracy by adjusting the continuous value $\lambda$ in the hypernetwork’s input weights. This results in a complete, continuous Pareto front approximation, making performance indicators like Cardinality indicators, which calculate the number of points in the Pareto front approximation, less suitable. Likewise, distribution and spread indicators, which calculate the dispersion among discrete points, have little practical significance.
Moreover, when discussing multi-objective optimization problems in the IR context, there is often no ground truth Pareto front for the problem, making it impossible to measure the proximity between the ground truth and the approximation. Finally, among the convergence and distribution indicators, only the hypervolume indicator [33] is more appropriate for the current discussion context, and its calculation is feasible. Below, we will introduce the hypervolume indicator.
The hypervolume indicator [33] is a widely used performance metric in multiobjective optimization that measures the volume of the objective space dominated by a Pareto front approximation relative to a reference point. This indicator provides a comprehensive assessment of both convergence and diversity of the Pareto front.
Mathematically, the hypervolume $\mathrm{HV}$ for a given Pareto front approximation $P$ and a reference point $r$ is defined as:
$\mathrm{HV}(P,r)=\\ \text{vol}\left(\bigcup_{x\in P}\left[f_{1}(x),r_{1}\right]\times\left[f_{2}(x),r_{2}\right]\times\cdots\times\left[f_{m}(x),r_{m}\right]\right)$ (1)
where $f_{i}(x)$ is the value of the $i$-th objective function for solution $x$.
The hypervolume indicator calculates the volume of the $m$-dimensional space covered by the union of hyperrectangles defined by each solution in $P$ and the reference point $r$. The higher the hypervolume value, the better the approximation of the Pareto front, indicating better convergence and spread.
The choice of the reference point $r$ is critical as it influences the hypervolume value. It is usually set to a point that is dominated by all solutions. The hypervolume indicator is a powerful tool for evaluating the quality of Pareto front approximations in multiobjective optimization, balancing both convergence to the true Pareto front and diversity of the solutions.
5.2 Dataset

### v2

@@ 4.1.4 Other Methods
There are also some control methods in machine learning, which is not strictly belong to the above three classifications. Some studies [85, 16, 15, 86] propose the concept of Disentanglement , which decouples user interests into specific dimensions within the latent space to facilitate control. Specifically, Multi-VAE and Supervised $\beta$-VAE [85, 15] provide the ‘knobs’ for users, with each knob corresponding to an item aspect. The latent space where generative factors (here, a preference towards an item category like genre) are captured independently in their respective dimensions, thereby enabling predictable manipulations. LP [16] models user-item interactions and user requirements in a linear space, enabling controllable user updates by integrating user requirements into user features through a weighted summation approach. CGIR [86] propose a weakly-supervised method that can learn a disentangled item representation from user-item interaction data and ground the semantic meaning of attributes to dimensions of the item representation. During inference, CGIR start from the reference item and “walk” along the direction of the modification in the item representation space to retrieve a sequence of items in a gradient manner. Test-Time Adaption (TTA) means the algorithms do not change the model through retraining manner. especially in the scenario of online recommendation, where time consuming is quite strict, some works focus on how to change the model to adapt the dynamic environment directly in test stage. Reinforcement Learning (RL) is a kind of classic method to achieve controllability. Since these algorithms learn from the interaction with the environment (actually the reward function of the environment), it is possible for us to control the algorithms through dedicatedly designing the reward function. LangPTune [17] leverages reinforcement learning to enable the encoder to better encode implicit user interaction sequences into interpretable text. Similarly, Lu et al. [5] adopt a reinforcement learning approach to design fine-grained reward functions for control rules, allowing the model to accurately respond to specific control instructions. Additionally, inspired by the exceptional general intelligence of Large Language Models (LLMs) [63], researchers have begun to explore the LLMs application in recommender systems and other fields. RecLM-gen [5] constructs Supervised Fine-Tuning (SFT) tasks, augmented with labels derived from a conventional recommender model, explicitly improving LLMs’ proficiency in adhering to recommendation specific instruction which could be regarded as task description $\bm{s}_{\mathrm{desc}}$ as defined in Sec.2. For example, DLCRec [87] is a novel framework designed to enhance diversity control in LLM-based recommender systems. By decomposing the recommendation task into three supervised fine-tuning tasks-genre prediction, genre filling, and item prediction—it enables fine-grained diversity control guided by user-defined signals. Apart from the aforementioned technical approaches, recent works have leveraged large language models (LLMs) to enable interpretable and controllable Natural Language Interest Editing (NLIE). Notable examples include LACE [6], TEARS [9], and LangPTune [17], among others.
Today, an increasing number of techniques are being applied to the field of controllable learning, and we look forward to the emergence of more novel and effective algorithms in the future.
@@
Table 1: Summary of representative controllable learning (CL) methods for information retrieval. “Who” denotes the entity responsible for control (see Section 3.2), including user-centric control and platform-mediated control. “What” refers to what is task target (see Section 3.1) including, multi-objective control, user portrait control, scenario adaptation control. “CL Tech.” describes the techniques for implementing control (see Section 4.1) including rule-based, pareto optimization, hypernetwork and others. “Where” indicates the position where control is applied in the testing stage (see Section 4.2) including pre-processing methods, in-processing methods, and post-processing methods.
@@
Method Year What Who CL Tech. Where
PadiRec [4] 2024 multi-objective control platform-mediated control hypernetwork in-processing
RecLM-gen [5] 2024 multi-objective control platform-mediated control SFT, RL in-processing
IFRQE [7] 2024 user portrait control user-centric control others pre-processing
TEARS [9] 2024 user portrait control user-centric control RL, NLIE in-processing
CMBR [10] 2024 user portrait control user-centric control SFT, RL in-processing
LangPTune [17] 2024 user portrait control user-centric control RL, NLIE in-processing
CCDF [18] 2024 multi-objective control platform-mediated control others in-processing
CMR [3] 2023 multi-objective control platform-mediated control hypernetwork in-processing
LACE [6] 2023 user portrait control user-centric control NLIE pre-processing
UCR [11] 2023 user portrait control user-centric control others pre-processing
Hamur [13] 2023 scenario adaptation control platform-mediated control hypernetwork, TTA in-processing
HyperBandit [12] 2023 scenario adaptation control platform-mediated control hypernetwork in-processing
PEPNet [14] 2023 scenario adaptation control user-centric control hypernetwork in-processing
SAMD [20] 2023 scenario adaptation control platform-mediated control hypernetwork in-processing
DTRN [21] 2023 scenario adaptation control user-centric control hypernetwork in-processing
MoFIR [82] 2022 multi-objective control user-centric control pareto optimization in-processing
UCRS [2] 2022 multi-objective control user-centric control others pre-processing
PAPERec [81] 2021 multi-objective control user-centric control pareto optimization in-processing
Supervised $\beta$-VAE [15] 2021 user portrait control user-centric control Disentanglement in-processing
ComiRec [1] 2020 multi-objective control platform-mediated control others post-processing
LP [16] 2020 user portrait control user-centric control Disentanglement in-processing
MMR [39] 1998 multi-objective control platform-mediated control rule-based post-processing
@@ 4.2 Where to Control in Information Retrieval Models
In this section, we categorize control learning methods into three distinct categories based on the inference process during the testing phase. Specifically, we classify them based on the stage at which the control function is applied: methods that adjust model inputs before inference are termed pre-processing methods, those that operate on model parameters during inference are termed in-processing methods, and those that act on model outputs after inference are termed post-processing methods.
@@ 4.2.1 Pre-Processing Methods
Pre-processing methods achieve the task target solely by transforming the model inputs without adjusting the model itself. Transformation of model inputs includes directly concatenating the task description onto the original feature vector. These methods are akin to in-context learning methods in large language models, where the task description serves as the prompt.
@@ 4.2.2 In-Processing Methods
In-processing methods adaptively adjust the parameters or, more loosely speaking, the hidden states of models upon receiving the task description and context to achieve the task target. The CMR framework [3] harnesses the input preference vector to guide the hypernetwork, which in turn generates network parameters tailored to the desired balance of objectives, thus achieving the desired control during the testing stage. HyperBandit [12] utilizes the periodic time information to inject to a hypernetwork, modeling the relationship between user preference with corresponding time block, achieving efficient user preference adaptation during testing stage. Similarly, the CCDF [18] employs a hyperparameter, denoted as ‘$k$’, to directly manipulate the number of categories presented in the top-$k$ recommendations, allowing for precise control over the diversity of the recommended content. These methods demonstrate the utility of in-processing strategies in enhancing the adaptability and performance of recommender systems.
@@ 4.2.3 Post-Processing Methods
Post-processing methods, including reranking and result diversification, are crucial for refining the output of recommender systems. ComiRec [1] leverages a final aggregation module that balances dual objectives—accuracy and diversity—through a weighted summation approach to determine the top-k recommendations. This method ensures that the final selection of recommendations is not only precise but also diverse, catering to a broader range of user interests. Some works achieve controllability by re-ranking. MMR [39] is a post-processing technique in recommender systems that aims to balance relevance and diversity in the recommended items. By iteratively selecting items that maximize a combined criterion of relevance to the user’s query and minimal redundancy with already selected items, it ensures that the final recommendation list is both pertinent and varied.
@@ 5.1.1 Single-Objective Metrics
The objectives pursued by users and platforms can vary significantly. Users tend to prioritize the accuracy and novelty of recommendations, while platforms place greater emphasis on long-term revenue, recommendation efficiency, and other operational goals. In this chapter, we present a range of commonly used metrics for single-objective recommendation, categorized from various perspectives.
Accuracy. Accuracy in recommender systems refers to how well the recommended items match the actual preferences or interests of users. It measures the system’s ability to provide relevant items that the user is likely to engage with, such as clicking, purchasing, or interacting. Common metrics include:
@@
NDCG [40] evaluates the effectiveness of information retrieval systems by accounting for both the relevance and rank of retrieved documents. It applies a logarithmic discount to lower-ranked items and normalizes against an ideal ranking for consistency. This makes NDCG essential for comparing search engine and recommender system performance across different queries:
$\mathrm{NDCG}@k$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathrm{DCG}_{i}@k}{\mathrm{IDCG}_{i}@k},$
$\mathrm{DCG}_{i}@k$ $=\sum_{j=1}^{k}\frac{2^{y_{ij}}-1}{\log_{2}(j+1)},$
where $N$ is the number of test samples, $y_{ij}\in\{0,1\}$ denotes the label of the $j$-th item.
@@
Precision is a metric used to evaluate the accuracy of information retrieval systems by measuring the proportion of retrieved documents that are relevant. It is calculated as the number of relevant documents retrieved divided by the total number of documents retrieved. High precision indicates that the system retrieves mostly relevant items, minimizing the presence of irrelevant information. This metric is particularly important in scenarios where presenting irrelevant items can negatively impact user experience, such as in search engines, recommender systems, and spam filtering:
$\mathrm{Precision}@k=\sum_{i=1}^{N}\frac{\widehat{L}^{k}_{i}\cap L^{k}_{i}}{\widehat{L}^{k}_{i}},$
where $\widehat{L}^{k}_{i}$ is the top-$k$ list outputted by the IR model, $L^{k}_{i}$ is the ground-truth top-$k$ list.
@@
Recall measures the effectiveness of information retrieval systems by determining the proportion of relevant documents that have been successfully retrieved from the total relevant documents available. It is calculated as the number of relevant documents retrieved divided by the total number of relevant documents. High recall indicates that the system retrieves most of the relevant items, which is essential in tasks where missing relevant information can have significant consequences, such as medical research or legal discovery:
$\mathrm{Recall}@k=\sum_{i=1}^{N}\frac{\widehat{L}^{k}_{i}\cap L^{k}_{i}}{L^{k}_{i}}.$
@@
Hit Rate evaluates the performance of recommender systems by assessing the presence of relevant items within the top-N recommendations provided to users. It is calculated as the number of users for whom at least one relevant item is included in the top-N recommendations, divided by the total number of users. Hit Rate is particularly useful for understanding the effectiveness of recommendation algorithms in scenarios where presenting at least one relevant option can significantly impact user satisfaction, such as in e-commerce or content streaming platforms:
$\mathrm{Hit}@k=\sum_{i=1}^{N}\mathbb{I}(\widehat{L}^{k}_{i}\cap L^{k}_{i}\neq\emptyset),$
where $\mathbb{I}(\cdot)$ is an indicator function.
Diversity. Diversity in recommendation systems refers to the variety of items recommended to users. A diverse recommendation list includes items that are different from each other, avoiding redundancy and ensuring a broad range of options. Common metrics include:
@@
$\alpha$-NDCG [41] extends the NDCG metric to evaluate the diversity in information retrieval systems. By incorporating a parameter $\alpha$, it penalizes redundancy and rewards the retrieval of diverse, relevant documents. This makes $\alpha$-NDCG particularly useful for tasks like web search and recommender systems, where presenting varied content is crucial:
$\alpha\text{-}\mathrm{NDCG}@k$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\alpha\text{-}\mathrm{DCG}_{i}@k}{\alpha\text{-}\mathrm{IDCG}_{i}@k},$
$\alpha\text{-}\mathrm{DCG}_{i}@k$ $=\sum_{j=1}^{k}\sum_{l=1}^{m}\frac{t_{j,l}(1-\alpha)^{c_{j,l}}}{\log_{2}(j+1)},$
where $m$ is the number of subtopic. $t_{j,l}=1$ if the $j$-th item covers subtopic $l$ and $t_{j,l}=0$ otherwise. $c_{j,l}$ is the number of times the subtopic $l$ being covered by items prior to the $j$-th item.
@@
ERR-IA [42] evaluates the effectiveness of information retrieval systems by considering user intent in the evaluation process. ERR-IA extends the Expected Reciprocal Rank (ERR) metric by incorporating a probabilistic approach to user intents, ensuring that the evaluation reflects the variety of user needs. It models user satisfaction as a function of the relevance of retrieved documents and their alignment with multiple user intents, making ERR-IA particularly useful for web search and other scenarios where understanding and catering to diverse user intents is crucial:
$\mathrm{ERR}\text{-}\mathrm{IA}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{k}\frac{1}{j}\sum_{l=1}^{m}\frac{1}{m}\frac{t_{il}}{2^{c_{j,l}}+1}.$
@@
$\mathrm{Coverage}@k=\frac{\cup_{i=1}^{N}\widehat{L}^{k}_{i}}{|\mathcal{I}|},$
where $\mathcal{I}$ is the set of all items.
Fairness. Fairness [90, 91] in recommender systems aims to ensure that all users, regardless of demographic or social group, are treated equitably in the recommendations they receive. Group fairness is a subfield that focuses on ensuring that different groups (e.g., gender, race, age) are not disadvantaged by the recommendation process. We listed several commonly used group fairness metrics.
Demographic Parity (DP) ensures that different groups receive similar rates of recommendations, regardless of their past interactions or preferences:
$\mathrm{DP}=\left|\frac{\sum_{i=1}^{|\mathcal{S}_{0}|}\hat{y}_{i}}{|\mathcal{S}_{0}|}-\frac{\sum_{i=1}^{|\mathcal{S}_{1}|}\hat{y}_{i}}{|\mathcal{S}_{1}|}\right|,$
where $\mathcal{S}_{0}$ and $\mathcal{S}_{1}$ represent different groups divided based on sensitive features such as gender, $\hat{y}_{i}$ denotes the score predicted by the recommendation model.
Equal Opportunity (EO) ensures that groups with same feedback (e.g., clicks, purchases) are equally likely to receive relevant recommendations:
$\mathrm{EO}=\sum_{y\in\{0,1\}}\left|\frac{\sum_{i=1}^{|\mathcal{S}_{0}^{y}|}\hat{y}_{i}}{|\mathcal{S}_{0}^{y}|}-\frac{\sum_{i=1}^{|\mathcal{S}_{1}^{y}|}\hat{y}_{i}}{|\mathcal{S}_{1}^{y}|}\right|,$
where $\mathcal{S}_{0}^{y}$ and $\mathcal{S}_{1}^{y}$ represent different groups divided based on sensitive features and ground-truth labels.
@@
$\mathrm{ISO}\text{-}\mathrm{index}=\lambda\cdot\mathrm{Diversity}+(1-\lambda)\cdot\mathrm{Fairness},$
where $\lambda$ is a hyper-parameter.
Novelty is a metric that favors recommendations of items that are different from those the user has already encountered. It emphasizes introducing fresh and less familiar content, rather than recommending similar or previously interacted items. Following [92], novelty is defined as the degree to which recommended items differ from the user’s past interactions, with a higher novelty value indicating more diverse and less repetitive recommendations:
$\mathrm{Novelty}@k=\sum_{i=1}^{N}\sum_{j=1}^{|\widehat{L}^{k}_{i}|}\frac{\log(\mathrm{Pop}(\hat{l}_{i,j})+1)}{|\widehat{L}^{k}_{i}|},$
where $\hat{l}_{i,j}$ is the $j$-th item in the ranking list $\widehat{L}^{k}_{i}$, $\mathrm{Pop}(\hat{l}_{i,j})$ denotes the popularity of item $\hat{l}_{i,j}$.
@@ 5.1.2 Multi-Objective Optimization Metrics
According to [43], performance indicators in multiobjective optimization are mainly focus on continuous Pareto front approximation. However, in the field of controllable multi-objective optimization, the Pareto front approximation is not obtained through evolutionary algorithms that yield a finite (discrete) set of points. For example, [1] achieves diversity control by adjusting the continuous value $\lambda$ in the aggregation module, and similarly, [3] controls diversity and accuracy by adjusting the continuous value $\lambda$ in the hypernetwork’s input weights. Thus, we present the evaluation metrics for discrete Pareto front approximations:
Hypervolume[93] measures the volume of the objective space dominated by a set of solutions $S$, bounded by a reference point $z$. It evaluates both convergence and diversity. A larger HV value indicates better performance:
$\mathrm{HV}=\lambda\left(\bigcup_{i=1}^{|S|}\prod_{j=1}^{m}\left[f_{j}^{(i)},z_{j}\right]\right),$
where $S$ is the solution set, $f_{j}^{(i)}$ is the objective value of solution $i$ in the $j$-th objective, $z_{j}$ is the reference point in the $j$-th objective, and $\lambda(\cdot)$ denotes the Lebesgue measure.
R2 [94] is a scalarization-based metric used to evaluate the quality of a solution set in multi-objective optimization. It measures the proximity of the solution set to an ideal set based on predefined weight vectors, without requiring the true Pareto front. This metric is defined as:
[…]
where $W$ is a set of weight vectors that represent user-defined preferences in the objective space, $f_{i}(x)$ is the value of the $i$-th objective function for a solution $x$, $P$ is the set of solutions being evaluated, and $m$ is the number of objectives.
In some studies, the continuous Pareto front can be obtained, allowing the measurement of the distance between the discrete solution set and the Pareto front.
Generational Distance [95] quantifies the convergence of the obtained solution set $P$ to the true Pareto front $P^{*}$. It computes the average Euclidean distance between each solution in $P$ and the nearest point in $P^{*}$:
$\mathrm{GD}=\left(\frac{1}{|P|}\sum_{i=1}^{|P|}d_{i}^{p}\right)^{\frac{1}{p}},$
where $|P|$ is the number of solutions in $P$, $d_{i}$ is the Euclidean distance from the $i$-th solution in $P$ to the nearest point in $P^{*}$, and $p$ is typically set to 2.
Inverted Generational Distance [96] evaluates both convergence and diversity by computing the average distance from each point in the true Pareto front $P^{*}$ to the nearest solution in the obtained set $P$:
$\mathrm{IGD}=\frac{1}{|P^{*}|}\sum_{j=1}^{|P^{*}|}d_{j},$
where $|P^{*}|$ is the number of points in $P^{*}$, and $d_{j}$ is the Euclidean distance from the $j$-th point in $P^{*}$ to the nearest solution in $P$.
5.2 Datasets